OpenStudy (anonymous):

Help with Substitution Method!!! Please show me how to solve these problems step by step.. 1. y=x-2 y=4x+1 2. y=x-4 y=-x+2 3. y=3x+1 y=5x-3 4. 2x-y=6 x+y=-3 5. 2x+y=8 y=x-7 6. 2x+3y=0 x+2y=-1 7. 3x-2y=7 x+3y=-5 8. -2x+y=0 5x+3y=11 9. 1/2x+1/3y=5 1/4x+y=10 Thanks!! :)))

OpenStudy (anonymous):

I will show the first one: find x: y=x-2 (take the 2 to the other side ) y+2=x now put x=y+2 into the second equation